OpenStudy (anonymous):

anyone can help me with fraction exponents?eg.x^1/4.I want to know how it is calculated manually...Please help

OpenStudy (kira_yamato):

the denominator of the exponent is n, and you take nth-root of the base.

jhonyy9 (jhonyy9):

exactly how have wrote before x on exponent 1/4 is equal 4rootx

OpenStudy (anonymous):

suppose you have 16^1/4 you can say it is same as (2^4)^1/4 = 2^(4*1/4) = 2^4/4 = 2^1 = 2